OpenStudy (anonymous):

The point(-11,2) is on the graph of y=f(x). Find a point on the graph of y=g(x), where g(x) = 1/2f(x)

OpenStudy (anonymous):

(-11,2) is on the graph of y=f(x) means that 2=f(-11). So g(x) = 1/2f(x) means g(-11) = 1/2 f(-11)=1/2*2=1. So (-11,1) is that point

OpenStudy (anonymous):

Great! Thank you for explaining. I got stuck at g(x) = -11

OpenStudy (anonymous):

g(x) = -11 ?

OpenStudy (anonymous):

As in g(-11) was what I ment.

OpenStudy (anonymous):

think of g(x) = 1/2 f(x) like this. g(x) is a function very simmilar to f(x). The only difference is that is alway half of the value of f(x), So for ex. when f(x) = 10, g(x) = 5, when f(x) = 2, g(x) = 1 and so on...

OpenStudy (anonymous):

so when f(-11) = 2 , then g(-11) = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (anonymous):

yw